OpenStudy (anonymous):

The baseball diamond is in the shape of a square with each side being 90 feet. If the catcher at home plate throws out a runner at second base, how far does he need to throw the ball?

OpenStudy (anonymous):

You're missing the hypotenuse, and to solve this problem, you need to apply Pythagorean Theorem. a^2 + b^2 = c^2

OpenStudy (anonymous):

a^2 + b^2 = c^2 90^2 + 90^2 = c^2 8100 + 8100 = c^2 16200 = c^2 The square root of 16200 is 127.28 feet.

OpenStudy (anonymous):

That is the answer :)